OpenStudy (anonymous):

The zero of f(x)=x-4 is..

Parth (parthkohli):

Zero is when f(x) = 0 Since f(x) = x - 4, we could equate \(x - 4 = 0\) Can you do it now?

OpenStudy (anonymous):

would it equal 4?

OpenStudy (callisto):

Yes.

OpenStudy (anonymous):

Ok thanks. What about: g(x)=2x+3

OpenStudy (callisto):

Find the zero again ?

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

i got the answer - (3/2)

OpenStudy (callisto):

Yes.

OpenStudy (anonymous):

i aint got that answer on my sheet tho. all i have is - 1/2 and -1 1/2

OpenStudy (callisto):

..... -(3/2) = - [(2+1)/2] = - 1 1/2

OpenStudy (anonymous):

Oh ok. I understand that.

OpenStudy (anonymous):

What about f(x)= (2x+1)/(3)

OpenStudy (anonymous):

i got -1/2

OpenStudy (callisto):

Yes.

OpenStudy (anonymous):

I think im getting the hang of this.

OpenStudy (callisto):

Hmm?